# Right Reward Right Time for Federated Learning

### v1

@@ 2.2 Learning Workflow
R3T is designed to perform model training, storage, and reward calculation and allocation in a transparent, fair, and auditable manner, leveraging blockchain and FL technologies. It integrates a time-aware contract theory-based mechanism to address right-time and right-reward allocation where CLPs phenomenon exists. The learning process of R3T involves the following steps. 1) Prepare contracts. The cloud server initiates its naive FL model ${w}$, designs a set of contract items, and distributes them to $N$ registered clients. 2) Sign contracts & Download global model. Each client signs a contract item to participate in the training process and downloads the global model ${w}$. 3) Local training. Clients train ${w}$ using their private data $\mathcal{D}$. 4) Upload updates. Clients send their trained model updates, metadata (e.g., IDs, addresses, chosen contracts), and hashed trained model updates to their associated edge servers after training [^3] We assume that clients are truthful in reporting their updates to the cloud. If required, the cloud can verify their updates using the Trusted Execution Environments [40] or reputation systems [41]. Besides, given the traceability property of the underlying blockchain [24] and incentive mechanism design (e.g., incentive compatibility) presented later in Section 3, clients are encouraged to truthfully report to benefit from joining and contributing.. 5) Forward & Validate updates. Edge servers relay trained model updates to the cloud server to prevent delay in global model aggregation process and model convergence rate, and only cross-validate hashed trained model updates and metadata to ensure data integrity and maintain consensus, storing relevant data on the blockchain for transparency and arbitration in case of disputes. 6) Receive & Aggregate model updates [^4] To avoid the straggler effect, we employ a synchronous learning update scheme, which has a provable convergence [42, 27].. The cloud server aggregates them using an aggregation algorithm such as FevAvg [1]. 7) Settlement. The process iterates until the global model converges. Upon completion, the final global model is transmitted back to the cloud server, which issues rewards based on the signed contract items. The rewards, calculated using blockchain smart contracts, are distributed to clients to compensate for their contributions (i.e., time and effort) to the learning task. Edge servers are also incentivized by receiving rewards for validating and storing transaction data.
2.3 Time Frame, Client Type, Contract, and Strategy
@@
1) Time Frame: The time frame is divided into $T$ training rounds [^5] In this work, training rounds and time slots are used interchangeably., denoted as $\mathcal{T}=\{1,...,t,...,T\}$, in which CLPs are vital duration in determining the final model performance.
@@
2) Client Type: $N$ clients are classified into a set $\mathcal{K}=\{1,...,k,...,K\}$ of $K$ distinct types according to their available time and system capabilities (i.e., computation and communication resources, and the quantity of data) due to data and system heterogeneity among them in real-world conditions. We denote the capability of a client $n$ in terms of time and system resources of a type-$k$ client ($1\leq k\leq K$) by ${\theta}_{n,k}$ or ${\theta}_{k}$ [^6] For simplicity, we interchangeably represent ${\theta}_{n,k}$ and ${\theta}_{k}$ for a type-$k$ client $n$’s capability. Similarly, the contract item $\phi_{k}$ can be denoted as $\{{e_{n,k},t_{n,k},R_{n,k}}\}$ or $\{{e_{k},t_{k},R_{k}}\}$. The same notation applies to the model parameter $w$, which is introduced later..
@@
It signifies that higher client types have greater training capabilities and a higher willingness to participate in early-available time, especially in CLPs, as well as provide a huge amount of data. The cloud is unaware of the exact client types, knowing only the probability of a client belonging to a certain type $k$. Due to privacy regulations, direct collection of client-specific information is restricted, making it challenging to predict individual client capabilities. Instead, only aggregate statistical insights can be obtained through market research and surveys conducted with client consent [43].
3) Cloud’s Contract: To motivate clients to join early, share resources, and conduct training, the cloud must offer transparent and fair contribution-reward bundles, compensating clients for their efforts. Recognizing the importance of CLPs, the cloud offers extra bounties to attract high-quality, resource-rich clients to participate actively. Consequently, each client’s gain is influenced by their chosen time of participation in the training process. The cloud has a contract set $\bm{\phi}=\{\phi_{k}\}_{k\in\mathcal{K}}$, which contains $K$ contract items corresponding to $K$ client types. The contract item $\phi_{k}=\{e_{k},t_{k},R_{k}\}$ specifies the relationship between type-$k$ client’s effort, joining time, and reward. Specifically, $e_{k}$ is defined as the number of data units required for each training round, and $t_{k}$ is the training time slot, which is either in CLPs or in non-CLPs. Also, $R_{k}$ is the corresponding reward for type-$k$ client paid by the cloud server for finishing the training round, and is given as follows
@@
where $r_{k}$ is the basic salary and $B_{k}$ is the bonus, which has been shown as an effective payment structure to motivate clients’ efforts toward collective goals (i.e., enhancing global model performance) [44]. Specifically, clients are assured to have a base compensation of salary $r$ upon joining the training process. Based on their individual types, the timing of their participation, and the level of effort they contribute, bonuses are awarded. These bonuses direct clients participating in CLPs to contribute more efforts during this period, as well as providing additional rewards for the highest-contributing clients [45]. Besides, inspired by the demand-price relationships in [46], we propose a function $h(t_{k})$ which is a time-aware bonus unit function (i.e., bonus-time relationship function), where the crucial role of CLPs is known by clients, and is presented as follows
@@
where $\vartheta$ represents the bonus unit coefficient. That function offers high-quality clients bonuses when they choose to join the model training process at an early stage, given the importance of CLPs.
4) Client’s Strategy Space: Each client decides whether to participate in the training process, (if yes) how much effort they should exert and which time slot to join. The choice of different time slots may result in varying bonuses. Particularly, participating during the CLPs yields higher payoffs for clients who contribute significant effort, as training the global model with huge effort during these periods is crucial for improving final model performance.
2.4 Utilities
According to (1), $R_{k}$ is dependent of $e_{k}$, $t_{k}$, and $r_{k}$, thus each contract item is rewritten by $\phi_{k}=\{e_{k},t_{k},r_{k}\}$.
1) Clients: The utility of type-$k$ client $n$ can be defined by
@@ 2.4 Utilities
where $\left(\frac{1}{2}\delta e_{k}^{2}+\beta r_{k}\right)$ is the total contribution cost. This includes $\frac{1}{2}\delta e^{2}_{k}$, which is a quadratic cost function with respect to client’s effort (e.g., sensing, collecting, and training of data or using cloud services for these tasks) according to [21], with $\delta>0$ being a quadratic unit effort cost that varies on different learning tasks (e.g., generating data samples or training data samples). For example, training ChatGPT-3 with 45TB of compressed plaintext before filtering and 570GB after filtering eighty using 80 V100 GPUs costed more than \$2 million or generating 100,000 data points using Amazon’s Mechanical Turk service will cost around \$70,000 [47]. Additionally, $\beta r_{k}$ represents the time cost associated with joining time $t_{k}$, considered as a linear function of the reward, where $\beta>1$ is the time cost coefficient [21, 48].
2) Edge Server: The edge server $l$’s utility is defined as $\mathcal{U}_{l}={R_{l}}-v(f_{l}),$ where $R_{l}$ is the gain for data mining paid by cloud, $v(f_{l})$ is the mining cost function, and $f_{l}$ is the edge server’s computing resource such as CPU cycle frequency.
3) Cloud server: The cloud’s utility is the difference between the gain (i.e., FL model performance) and costs (i.e., rewards for clients’ contributions such as joining time and efforts in training the FL model, and for edge server in mining and validating data transactions) as follows:
@@
where $\lambda$ is a parameter adjusted by the cloud server to show its concerns for FL model performance (i.e., larger $\lambda$) or reward budget (i.e., smaller $\lambda$), $\bm{e}=\{e_{1},\dots,e_{K}\}$, $\bm{t}=\{t_{1},\dots,t_{K}\}$, and $\bm{r}=\{r_{1},\dots,r_{K}\}$. Without the loss of generality, $g(\cdot)$ is defined as a concave function [14, 49] with respect to the amount of effort and joining time. Specifically, with the same effort, each client can make different contributions and impacts on FL model performance at different training times.
2.5 Blockchain smart contract-based R3T procedure
Our focus is on the incentive mechanism design given the existence of CLPs in FL, which can be considered as a service and is built on top of the underlying blockchain [23, 24, 50], thereby issues of blockchain security and scalability are out of scope in this article. Specifically, R3T uses smart contracts to credibly enable client and cloud registration, effort record, data storage, and reward calculation and distribution without involvement of third parties. In case of disputes between clients and the cloud, the immutably recorded results on the blockchain can serve as reliable evidence for arbitration.
@@ 2.5 Blockchain smart contract-based R3T procedure
Fig. 3 illustrates the functionality of blockchain and smart contracts in the R3T, including four primary operations. 1) First, clients and the cloud server initiate their participation. During this process, the system records relevant information, such as IDs, and wallet addresses, for both clients and the cloud server to ensure that only registered participants can interact within the system. Additionally, the cloud server publishes a contract set $\phi^{*}$, where $\phi^{*}$ is a set of optimal contract items determined by the cloud server in Section 3, directly onto the blockchain for transparency and accessibility to clients. 2) Upon completing the training process, clients upload their strategy profiles or selected contract items $\phi_{k}^{*}$ with their IDs, the location of the model weights, and hashed model weights to the blockchain by invoking contributitonSubmit() function in the Contribution smart contract. 3) To enhance the scalability, minimize on-chain storage overhead, and provide robustness to single-point-of-failure attacks for the cloud server of R3T, dataStore() in the Contribution smart contract is employed. Global model weights are uploaded to private IPFS built by authorized edge servers while the resulting hash and corresponding training round number are recorded on-chain, ensuring data integrity and efficient and decentralized model management. 4) Finally, contract-based rewards $R_{k}$ (e.g., tokens or cryptocurrencies such as Ether in the Ethereum blockchain platform) are calculated and distributed through reward() function in the Reward smart contract.
In the next section, we analyze R3T under two information conditions over CLPs and non-CLPs. 1) Complete Information Case: Clients’ types are public and known to the cloud server and other clients. However, this case is impractical due to privacy leakage risks from exposing individual properties. 2) Incomplete Information Case: Clients’ types are private and only measured locally, but the cloud server may statistically know the distribution of clients’ types through historical market research or surveys [43].
3 R3T - The Proposed Incentive Mechanism for critical learning periods
3.1 Problem Formulation
@@
Each client is willing to join and contribute its effort to the training process, which guarantees IR and IC constraints and leads to the best utility for itself [21]. For example, realizing the importance of CLPs with a correspondingly huge amount of rewards announced by the cloud server [^7] That is considered as common knowledge [21]., each client tends to join and allocate efforts in this period.
@@
To maximize the utility, the cloud server offers the optimal contract set $\bm{\phi^{*}}=\{\bm{e^{*}},\bm{t^{*}},\bm{r^{*}}\}$ to clients. The optimal contract set is the solution to the following problem [^8] The total blockchain mining rewards are constant because the computing resources required for this task are the same across all edge servers in each global training round [35]. Thus, this factor does not affect the optimization and analysis in this paper, so we omit it here.:
@@
  • Objectives: each type-$k$ client $n$ aims to maximize its utility $\mathcal{U}_{n}(e_{k},t_{k},r_{k})$ expressed in (2.4).
Once $t_{k}$ and $r_{k}$ are given, we determine that the utility of a client in (2.4) is a strictly concave function with respect to its effort $e_{k}$. Hence, the optimal choice of effort $e_{k}$ for a type-$k$ client can be obtained by setting the first-order derivative of the client’s utility function with respect to $e_{k}$. Specifically,
@@
Substituting $e^{*}_{k}$ into (1), (2.4), (6), the utilities of a type-$k$ client and the cloud server are rewritten respectively as
@@
3.2 R3T Incentivization under Complete Information Case
@@
It is important to note that a larger $h(t_{k})$ means a smaller $t_{k}$ that is at the early training stage. The optimal contract design in (16a) is rewritten by
@@
3.3 R3T Incentivization under Incomplete Information Case
We now introduce the contract game in an incomplete information setting. Clients choose contract items (i.e., effort, joining time, and corresponding reward) offered by the cloud server. While the cloud server aims to optimize its utility (e.g., global model performance), clients seek to maximize their effort-based incentives.

### v2

@@ II-B Learning Workflow
R3T is designed to perform model training, storage, and reward calculation and allocation in a fair and auditable manner, leveraging blockchain and FL technologies. It integrates a time-aware contract theory-based mechanism to address right-time and right-reward allocation where CLPs phenomenon exists. The R3T’s one-training-round process involves the following steps. 1) Prepare contracts. The cloud initiates a global model w{w}, designs a set of contract items, and distributes them to $N$ registered clients. 2) Sign contracts & Download global model. Each client signs a contract item to participate in the training process and downloads the global model w{w}. 3) Local training. Clients train w{w} using their private data $\mathcal{D}$. 4) Upload updates. Clients send their trained model updates, metadata (e.g., IDs, addresses, chosen contracts), and hashed trained model updates to their associated edge servers after training [^3] We assume that clients are truthful in reporting their updates to the cloud. If required, the cloud can verify their updates using the Trusted Execution Environments [zhang2020enabling] or reputation systems [yu2024mathcal]. Besides, given the traceability property of the underlying blockchain [yu2023ironforge] and incentive mechanism design (e.g., incentive compatibility) presented later in Section III, clients are encouraged to truthfully report to benefit from joining and contributing.. 5) Forward & Validate updates. Edges relay trained model updates to the cloud to minimize latency in the model aggregation process, and cross-validate hashed trained model updates and metadata to ensure data integrity and maintain consensus, storing records on the blockchain for transparency and dispute arbitration. 6) Receive & Aggregate model updates [^4] To avoid the straggler effect, we employ a synchronous learning update scheme, which has a provable convergence [chen2019round, kairouz2021advances].. The cloud aggregates them using an aggregation algorithm such as FedAvg [mcmahan2017communication]. 7) Settlement. The process iterates until the global model converges. Upon completion, smart contracts calculate and distribute rewards to clients to compensate for their contributions (i.e., time and effort). Edge servers also receive rewards for validating transactions and storing data.
II-C Time Frame, Client Type, Contract, and Strategy
@@
1) Time Frame: The time frame is divided into $T$ training rounds [^5] In this work, training rounds and time slots are used interchangeably., denoted as $\mathcal{T}=\{1,...,t,...,T\}$, in which CLPs $\subset\mathcal{T}$ are vital duration in determining the final model performance.
@@
2) Client Type: $N$ clients are classified into a set $\mathcal{K}=\{1,...,k,...,K\}$ of $K$ distinct types according to their available time and system capabilities (e.g., computation resources, and the quantity and diversity of data) due to statistical and system heterogeneity among them in real-world conditions. We denote the capability of a client $n$ in terms of time and system resources of a type-$k$ client ($1\leq k\leq K$) by ${\theta}_{n,k}$ or ${\theta}_{k}$ [^6] For simplicity, we interchangeably represent ${\theta}_{n,k}$ and ${\theta}_{k}$ for a type-$k$ client $n$’s capability. Similarly, the contract item $\phi_{k}$ can be denoted as $\{{e_{n,k},t_{n,k},R_{n,k}}\}$ or $\{{e_{k},t_{k},R_{k}}\}$. The same notation applies to the model parameter $w$, which is introduced later..
@@
It signifies that higher client types have greater training capabilities and a higher willingness to participate early, especially in CLPs, as well as be able to contribute high efforts (e.g., a huge amount of closer-distribution-to-the-learning-task data [10557146]). The cloud is unaware of the exact client types, knowing only the probability of a client belonging to a certain type $k$. Due to privacy regulations, direct collection of client-specific information is restricted, making it challenging to predict individual client capabilities. Instead, aggregate statistical insights can be obtained through market research and surveys conducted with client consent [10.1145/3209582.3209588].
3) Cloud’s Contract: To motivate clients to join early, share resources, and conduct training, the cloud must offer transparent and fair contribution-reward bundles, compensating clients for their efforts. Recognizing the importance of CLPs, the cloud offers extra bounties to attract high-quality, resource-rich clients to participate actively. Consequently, each client’s gain is influenced by their chosen time of participation in the training process. The cloud has a contract set $\bm{\phi}=\{\phi_{k}\}_{k\in\mathcal{K}}$, which contains $K$ contract items corresponding to $K$ client types per training round. The contract item $\phi_{k}=\{e_{k},t_{k},R_{k}\}$ specifies the relationship between type-$k$ client’s effort, joining time, and reward. Specifically, $e_{k}$ is defined as the normalized local training effort, which can be determined jointly by the volume and the distribution of local data [10557146], and $t_{k}$ is the training time slot, which is either in CLPs or in non-CLPs. Also, $R_{k}$ is the corresponding reward for type-$k$ client paid by the cloud server for finishing the training round, and is given as follows:
@@
where $r_{k}$ is the basic salary and $B_{k}$ is the bonus, which has been shown as an effective payment structure to motivate clients’ efforts toward collective goals (i.e., enhancing global model performance) [joseph1998role]. Specifically, clients are assured to have a base compensation of salary $r$ upon joining the training process. Based on their individual types, the timing of their participation, and the level of effort they contribute, bonuses are awarded. These bonuses direct clients participating in CLPs to contribute more efforts during this period, as well as providing additional rewards for the highest-contributing clients [churchill1993sales]. Besides, inspired by the demand-price relationships in [SELCUK20191191], we propose a function $h(t_{k})$ which is a time-aware bonus unit function (i.e., bonus-time relationship function), where the crucial role of CLPs is known by clients, and is presented as follows:
@@
where $\vartheta$ represents the bonus unit coefficient. That function offers high-quality clients bonuses when they choose to join the model training process at an early stage, given the importance of CLPs. To identify CLPs, the relative change in the weighted-average training losses aggregated over all selected type-$k$ clients between two consecutive rounds can be monitored. If this loss reduction exceeds a pre-defined threshold, the current round is in CLPs [yan2023criticalfl].
4) Client’s Strategy Space: Each client decides whether to participate in the training process (if yes), how much effort they should exert, and which time slot to join. The choice of different time slots may result in varying bonuses. Particularly, participating during the CLPs yields higher payoffs for clients who contribute significant effort, as training the global model with huge effort during these periods is crucial for improving final model performance.
II-D Utilities
According to (1), $R_{k}$ is dependent on $e_{k}$, $t_{k}$, and $r_{k}$, thus each contract item is rewritten by $\phi_{k}=\{e_{k},t_{k},r_{k}\}$. We define the utilities of each client, edge server, and cloud server per training round as follows.
1) Clients: The utility of type-$k$ client $n$ can be defined by:
@@ II-D Utilities
where $\left(\frac{1}{2}\delta e_{k}^{2}+\beta r_{k}\right)$ is the total contribution cost. This includes $\frac{1}{2}\delta e^{2}_{k}$, which is a quadratic cost function with respect to client’s effort (e.g., sensing, collecting, and training of data or using cloud services for these tasks) according to [bolton2004contract], with $\delta>0$ being a quadratic unit effort cost that varies on different learning tasks (e.g., generating data samples or training data samples). For example, training ChatGPT-3 with 45TB of compressed plaintext before filtering and 570GB after filtering using 80 V100 GPUs costs more than \$2 million, or generating 100,000 data points using Amazon’s Mechanical Turk service will cost around \$70,000 [brown2020language]. Additionally, $\beta r_{k}$ represents the time cost associated with joining time $t_{k}$, considered as a linear function of the reward, where $\beta>1$ is the time cost coefficient [bolton2004contract, wang2023connectivity].
2) Edge Server: The edge server $l$’s utility is defined as $\mathcal{U}_{l}={R_{l}}-v(f_{l}),$ where $R_{l}$ is the gain from validating data transactions, $v(f_{l})$ is the validation cost function, and $f_{l}$ is the edge server’s computing resource (e.g., CPU-cycle frequency).
3) Cloud server: The cloud’s utility is the difference between the gain (i.e., FL model performance) and costs (i.e., rewards for clients’ contributions such as joining time and efforts in training the FL model, and for edge servers in validating data transactions) as follows:
@@
where $\lambda$ is a parameter adjusted by the cloud server to show its concerns for FL model performance (i.e., larger $\lambda$) or reward budget (i.e., smaller $\lambda$), $\bm{e}=\{e_{1},\dots,e_{K}\}$, $\bm{t}=\{t_{1},\dots,t_{K}\}$, and $\bm{r}=\{r_{1},\dots,r_{K}\}$. In each training round, multiple clients may select the same contract item $k$, and $R_{k}$ in (6) denotes the aggregate reward cost incurred by the cloud for all participating clients choosing item $k$ in that round; this aggregate notation keeps (6) compact while remaining consistent with the budget feasibility constraint, which is defined in the next section. Without the loss of generality, $g(\cdot)$ is defined as a concave function [kang2019incentive, zhou2019computation] with respect to the amount of effort and joining time. Specifically, with the same effort, each client can make different contributions and impacts on FL model performance at different training times.
II-E Blockchain smart contract-based R3T procedure
Our focus is on the incentive mechanism design given the existence of CLPs in FL, which can be considered as a service and is built on top of the underlying blockchain [wang2022infedge, yu2023ironforge], thereby issues of blockchain security and scalability are out of scope in this article. Specifically, R3T uses smart contracts to credibly enable client and cloud registration, effort record, data storage, and reward calculation and distribution without involvement of third parties. In case of disputes between clients and the cloud, the immutably recorded results on the blockchain can serve as reliable evidence for arbitration.
@@ II-E Blockchain smart contract-based R3T procedure
Fig. 3 illustrates the functionality of blockchain and smart contracts, including four main operations. 1) First, clients and the cloud initiate their participation. During this process, R3T records information, such as IDs, and wallet addresses, for both clients and the cloud server to ensure that only registered participants can interact within the system. Additionally, the cloud publishes a contract set $\phi^{*}$, where $\phi^{*}$ is a set of optimal contract items determined by the cloud in Section III, directly onto the blockchain for transparency and accessibility to clients. 2) Upon completing the training process, clients upload their strategy profiles or selected contract items $\phi_{k}^{*}$ with their IDs, the location of the model weights, and hashed model weights to the blockchain by invoking contributionSubmit() function in the Contribution smart contract. 3) To enhance the scalability, minimize on-chain storage overhead, and provide robustness to single-point-of-failure attacks for the cloud server of R3T, dataStore() in the Contribution smart contract is employed. Global model weights are uploaded to private IPFS built by authorized edge servers while the resulting hash and corresponding training round number are recorded on-chain, ensuring data integrity and efficient and decentralized model management. 4) Finally, contract-based rewards $R_{k}$ (e.g., tokens or cryptocurrencies such as Ether in the Ethereum blockchain platform) are calculated and distributed through reward() function in the Reward smart contract.
Next, we analyze R3T under two information conditions over CLPs and non-CLPs. 1) Complete Information Case: Clients’ types are public and known to the cloud and other clients. However, this case is impractical due to privacy leakage risks from exposing individual properties. 2) Incomplete Information Case: Clients’ types are private and measured locally, but the cloud may statistically know the distribution of clients’ types via market research or surveys [10.1145/3209582.3209588].
III R3T - The Proposed Incentive Mechanism for critical learning periods
III-A Problem Formulation
@@
Each client is willing to join and contribute its effort to the training process, which guarantees IR and IC constraints and leads to the best utility for itself [bolton2004contract]. For example, realizing the importance of CLPs with a correspondingly huge amount of rewards announced by the cloud server [^7] That is considered as common knowledge [bolton2004contract]., each client tends to join and allocate efforts in this period.
@@
To maximize the utility, the cloud server offers the optimal contract set $\bm{\phi^{*}}=\{\bm{e^{*}},\bm{t^{*}},\bm{r^{*}}\}$ to clients. The optimal contract set is the solution to the following problem [^8] The total blockchain mining rewards are constant because the computing resources required for this task are the same across all edge servers in each global training round [wang2023incentive]. Thus, this factor does not affect the optimization and analysis in this paper, so we omit it here.:
@@
  • Objectives: each type-$k$ client $n$ aims to maximize its utility $\mathcal{U}_{n}(e_{k},t_{k},r_{k})$ expressed in (II-D).
Once $t_{k}$ and $r_{k}$ are given, we determine that the utility of a client in (II-D) is a strictly concave function with respect to its effort $e_{k}$. Hence, the optimal choice of effort $e_{k}$ for a type-$k$ client can be obtained by setting the first-order derivative of the client’s utility function with respect to $e_{k}$. Specifically,
@@
Substituting $e^{*}_{k}$ into (1), (II-D), (6), the utilities of a type-$k$ client and the cloud server are rewritten respectively as
@@
III-B R3T Incentivization under Complete Case
@@
It is important to note that a larger $h(t_{k})$ means a smaller $t_{k}$ that is at the early training stage. The optimal contract design in (16a) is rewritten by:
@@
III-C R3T Incentivization under Incomplete Information Case
We now introduce the contract game in an incomplete information setting. Clients choose contract items (i.e., effort, joining time, and corresponding reward) offered by the cloud. While the cloud aims to optimize its utility (e.g., global model performance), clients seek to maximize their incentives.